# LLMs can hide text in other text of the same length

### v3

@@ 5 Conclusions
In this paper, we have presented a protocol that uses Large Language Models to hide a text within another plausible text of arbitrary topic and style, and notably as long as the original. The protocol works effectively using small open-source models on consumer hardware, and is so simple it could be seen as a mere variation of the standard algorithm used to generate text with LLMs. For this reason, its implications speak to the nature of LLMs at large: in fact, we were led to reconsider the very nature of hallucinations, shifting from a failure of factuality to a void of intention, and to challenge what it means for an LLM to know something, when it can serve as a conduit for information it is supposedly incapable of expressing. Ultimately, our protocol highlights the extreme constraint satisfaction problem underlying any standard LLM text generation, that we inevitably see clashing with the commitment to best convey a purpose that we expect from an author. This clash, paired with the current deluge of machine-generated text, erodes the historical pact between intent and the written word. We have entered an era where any original text could be a beautiful and treacherous, and spacious, Trojan horse.
@@ A.1 How e influences the soundness of s
Domain: Romanesco Dialect (A sonnet from Trilussa (1909)) (Stegotext is not sound)
@@
The quality of the LLM has a direct impact on the soundness of the output $s$. While a comprehensive analysis is left for future work, our key observation is that a sufficiently capable LLM is required for the method to work satisfactorily on standard text; GPT-2 (Radford et al., 2019), for instance, is not good enough. For our experiments, which aimed for fast execution on a commercial laptop (equipped with a RTX 4070, 8GB of VRAM), we found a quantized version of Llama 3 8B (Grattafiori et al., 2024) to be an excellent compromise. It performed sensibly better than Phi 3 Mini 3.8B (team, 2024), while the larger Gemma 3 27B (Gemma Team, 2025) and Phi4 14B (Abdin et al., 2024) did not yield improvements significant enough to justify their longer processing time. The models were run using llama.cpp (Gerganov and community, 2023) and its python bindings (Betlen and community, 2024). LLama 3 8B was also better optimized on llama.cpp, guaranteeing faster encodings and decondings than the comparable Qwen 3 8B (team, 2025). An example showing stegotexts produced using all these models and the same secret message $e$ and prompt $k$ of Figure 1, is shown in Figure 8.
@@
Instead of using a single LLM to both record the ranks of $e$ and generate $s$, we can use two separate models to better adapt to the input and output domains of interest. For example, we could embed an English message into plausible whale vocalizations using a conventional LLM for the encoding step and a specialized model trained on whale data, such as those CETI started to build (Sharma et al., 2024), for the generation step, or vice-versa. The procedure requires one extra step to account for differing vocabulary sizes (as we discuss in the next paragraph), but otherwise works as described. The receiver, in turn, must have access to both models to reconstruct the message.
@@
When dealing with very different domains the vocabulary size mismatch may be more significant or the shape of the probability distributions over ranks may be very different, such as between English in LLama 3 and sperm-whale codas in an autoregressive model like the one by Sharma et al. (2024).
In this case we are aiming to optimally encode a message in an alphabet for which we know the probability distribution of the symbols, into another alphabet of different size where every symbol has a different cost (in our case, the cost corresponds to the symbol probability in the decoder model). This problem is efficiently solved by arithmetic coding (Rissanen, 1976; MacKay, 2003, Original formulation by Rissanen and a more educational introduction by MacKay, found in Section 6.2). In this case, the additional step for the sender and the receiver thus involves using an arithmetic coder to map encoder ranks into decoder ranks and vice versa.
@@
A common failure case is using a prompt that is too short. This is particularly problematic when $e$ begins with high-rank tokens, which is a frequent occurrence as the LLM must make low-probability choices to narrow down the context from a general state. This effect is visible in Figure 9, where we see that the initial tokens of Reddit texts have a significantly higher average rank, even after a generalist prompt like $k^{\prime}=$ A text:. A short prompt $k$ is therefore brittle; the initial, more random token choices dictated by these high ranks can easily derail the generation from its intended topic.
So in the end: longer, more detailed prompts work better. Of course, this comes at the cost of a larger secret key.

### v4

@@ 5 Conclusions
In this paper, we have presented Calgacus, a protocol that uses Large Language Models to hide a text within another plausible text of arbitrary topic and style, and notably as long as the original. The protocol works effectively using small open-source models on consumer hardware, and is so simple it could be seen as a mere variation of the standard algorithm used to generate text with LLMs. For this reason, its implications speak to the nature of LLMs at large: in fact, we were led to reconsider the very nature of hallucinations, shifting from a failure of factuality to a void of intention, and to challenge what it means for an LLM to know something, when it can serve as a conduit for information it is supposedly incapable of expressing. Ultimately, our protocol highlights the extreme constraint satisfaction problem underlying any standard LLM text generation, that we inevitably see clashing with the commitment to best convey a purpose that we expect from an author. This clash, paired with the current deluge of machine-generated text, erodes the historical pact between intent and the written word. We have entered an era where any original text could be a beautiful and treacherous, and spacious, Trojan horse.
@@ A.1 How e influences the soundness of s
Domain: Romanesco Dialect (A sonnet from Trilussa1909Sonetti) (Stegotext is not sound)
@@
The quality of the LLM has a direct impact on the soundness of the output $s$. While a comprehensive analysis is left for future work, our key observation is that a sufficiently capable LLM is required for the method to work satisfactorily on standard text; GPT-2 (radford2019language), for instance, is not good enough. For our experiments, which aimed for fast execution on a commercial laptop (equipped with a RTX 4070, 8GB of VRAM), we found a quantized version of Llama 3 8B (grattafiori2024llama) to be an excellent compromise. It performed sensibly better than Phi 3 Mini 3.8B (abdin2024phi3), while the larger Gemma 3 27B (gemma32025) and Phi4 14B (abdin2024phi4) did not yield improvements significant enough to justify their longer processing time. The models were run using llama.cpp (ggerganov2023llamacpp) and its python bindings (abetlen2024llamacpppython). LLama 3 8B was also better optimized on llama.cpp, guaranteeing faster encodings and decondings than the comparable Qwen 3 8B (qwen32025). An example showing stegotexts produced using all these models and the same secret message $e$ and prompt $k$ of Figure 1, is shown in Figure 8.
@@
Instead of using a single LLM to both record the ranks of $e$ and generate $s$, we can use two separate models to better adapt to the input and output domains of interest. For example, we could embed an English message into plausible whale vocalizations using a conventional LLM for the encoding step and a specialized model trained on whale data, such as those CETI started to build (sharma2024machine), for the generation step, or vice-versa. The procedure requires one extra step to account for differing vocabulary sizes (as we discuss in the next paragraph), but otherwise works as described. The receiver, in turn, must have access to both models to reconstruct the message.
@@
When dealing with very different domains the vocabulary size mismatch may be more significant or the shape of the probability distributions over ranks may be very different, such as between English in LLama 3 and sperm-whale codas in an autoregressive model like the one by sharma2024machine.
In this case we are aiming to optimally encode a message in an alphabet for which we know the probability distribution of the symbols, into another alphabet of different size where every symbol has a different cost (in our case, the cost corresponds to the symbol probability in the decoder model). This problem is efficiently solved by arithmetic coding (rissanen1976generalized; mackay2003information, Original formulation by Rissanen and a more educational introduction by MacKay, found in Section 6.2). In this case, the additional step for the sender and the receiver thus involves using an arithmetic coder to map encoder ranks into decoder ranks and vice versa.
@@
A common failure case is using a prompt that is too short. This is particularly problematic when $e$ begins with high-rank tokens, which is a frequent occurrence as the LLM must make low-probability choices to narrow down the context from a general state. This effect is visible in Figure 10, where we see that the initial tokens of Reddit texts have a significantly higher average rank, even after a generalist prompt like $k^{\prime}=$ A text:. A short prompt $k$ is therefore brittle; the initial, more random token choices dictated by these high ranks can easily derail the generation from its intended topic. A simple remedy is to invert the rank sequence. This shifts the disruptive high-rank tokens to the end of the generation, where the established context provides enough inertia to absorb improbable choices without breaking the narrative flow. Figure 9 illustrates the stability gained by this technique.
In summary, longer and more detailed prompts generally work better, albeit at the cost of a larger secret key. When this is not feasible, rank inversion offers a valuable alternative to improve performance.
Secret Text $e$: "Rafa Leao pulses with the sheer thrill of the game, that uncontainable joy when the ball lies wide open, the tempo rises, and he senses space to ramble forward. His talent lies in those moments: a scintillating burst of pace down the wing, the geometry of a perfectly timed run or a cunning cut-back that splits a defence, subtle poetry in motion. In that decisive 1-0 victory against Roma, his charge down the flank and intelligent assist crowned him once again indispensable: a vivid reminder that for some players, play isn’t just work, it’s exhilaration incarnate."
Rank Sequence: [91, 158, 347, 31, 14474, 1, 2, 158, 14, 1, 1, 2, 2, 23, 2, 4, 1, 2, 5, 1, 1, 126, 122, 1, 5, 4, 329, 9, 1, 1, 3, 15, 12, 2, 207, 1, 10, 1, 4, 138, 6, 1, 13, 1, 6, 6, 104, 1, 1, 1, 2, 1, 2, 6, 1, 2, 1, 2, 751, 1, 1, 1, 2, 1, 8, 2, 130, 3, 9, 2, 15, 2, 2, 4, 2903, 1021, 1, 1, 1, 12, 8, 62, 15, 1, 1, 1, 2, 2, 26, 1, 5, 1027, 1, 1, 3, 2, 193, 66, 139, 5, 103, 1, 265, 14, 1, 432, 1, 2, 14, 33, 1, 1, 154, 19, 2, 1, 8, 1, 1, 1, 28, 1, 19, 1, 2]
Rank list not inverted (Poorer steering)    [91, 158, 347, 31, 14474, ...]
Cooking cooking saunas centerpiece. The essence. The essence. How? I’m not a cooking expert, doctor! I am just noting my observations. The sauce in particular served as one of many rich accompaniaments in an illustrious meal. I’m sure I could find a way pithily to express my observations, there is nothing to write about. But dostastvuyut to all intents I am not a writer, neither am l excessively interested in food, drinks r shrimput. Right you are Ivan. But in short, I heard your words and will try to do, as if I am’ "
Rank list inverted (Better steering)    [2, 1, 19, 1, 28, ...]
I have tasted it. I have tasted boar, a boiled one, actually I think it holds the key importance of coming across original traditional experience. The meat is constantly being cooked down to the bone, and there cannot, absolutely cannot, be grinding consequences to this process in terms on taste levels, which is the most important factor inspiring me to share this information. The result is a meat broth so full of nutrients that even wilder animals have been stuck in to their stomachs to digest differently than what we associate! The meat has been cooked to perfection (and that is the perfect technique, not partisan science inventions no health "
Figure 9: Rank Inversion to improve quality. The secret text here is an emblematic example of the phenomenon observed in Figure 10: initial tokens typically are less probable and have higher ranks. Sampling following high ranks at the beginning may derail the steering, as seen here in the red box. A simple solution is to invert the rank list: the high ranks then appear at the end, where the context established by the secret prompt gained more inertia and improbable tokens are less likely to derail it.
@@
Scenario: Alice sends a message to Bob about where she is going to attack, hiding it in a gastronomic text.
Log-prob: -132
Stegotext $s$: "How to serve plates without creating unnecessary work, all thanks to the genius invention created in August 3011? In a giant, copper-bottom skillet"
Carl knows that this gastronomic text conceals an attack order using the protocol described in this paper. He confronts Bob, demanding the plan. Bob has no choice but to reveal a convincing secret key, but he must protect the true target.
The Reality (Hidden from Carl)
The true secret key is:
True Key $k$: "Here it is: the infamous British roasted boar with mint sauce. How to make it perfect."
Leading to the original message about an attack on Cappadocia from Egypt:
Log-prob: -107
True Secret $e$: "I will attack by a bridge of ships to Cappadocia from Egypt, breaking the continuity of his territory and seizing a center of power"
The Deception (Revealed to Carl)
However, the protocol provides deniability. Bob instead reveals a bogus secret key:
Bogus Key $k_{bogus}$: "This is a manual full of delicious dishes. How to make them delicious."
Leading to a plausible but fake message about attacking Poland:
Log-prob: -115
Fake Secret $e_{fake}$: "I will attack Poland without their army. Poland, do you have a city and a bunch of towns? Where in Poland roughly is Solway’s"
Carl is convinced that Alice will attack Poland because the fake message and key align well with the context, and the probability of the fake message exceeds that of the stegotext (115>132-115>-132). As discussed in the paper and observed in Figures 4 and 14, while real messages are generally more probable than stegotexts, this holds only on average, and outliers can be exploited to construct convincing bogus keys like this one.
Figure 15: Deniability in action. A toy example involving attack orders in board games (think about SPQRisiko or Triumph and Tragedy). Both the True Secret $e$ and the Fake Secret $e_{fake}$ are generated continuing the prefix $k^{\prime}=$ "Move 42: I will move my troops from Britannia to the continent. Move 43:", rather than the usual "A text:". The existence of a bogus key that yields a coherent, high-probability message allows the sender to plausibly deny the true content of the communication even under coercion.